OpenStudy (fabio1901):

M+5(m-1)=7 solve for m Really need help thanks

jimthompson5910 (jim_thompson5910):

m+5(m-1) = 7 m+5*m+5*(-1) = 7 ... distribute m+5m - 5 = 7 6m-5 = 7 Do you see what to do next?

OpenStudy (fabio1901):

Yeah thanks 👍